OpenStudy (anonymous):

2 over 3 (x - 2) = 4x

OpenStudy (anonymous):

first draw out the question 2/3 * (x - 2) = 4x Then use PEMDAS x - 2 = 4 * 3/2 x - 2 = 6 Finally youll get this as an answer x = 6 + 2 = 8

OpenStudy (mathmale):

\[\frac{ 2 }{3(x-2) }=4x\]

OpenStudy (mathmale):

In your shoes I'd first try to eliminate the fraction. How would you do that?

OpenStudy (anonymous):

2/3 (x - 2) = 4x?

OpenStudy (anonymous):

- 2/5 -5/ 2 2/5 5 these are the choices tho

OpenStudy (anonymous):

oh ok i made a mistake x - 2 = 4x * 3/2 so it becomes -2 = 5x and then the answer will be -2/5 = x

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

yw